Prestack exploding reflector modeling: The crosstalk problem

Claudio Guerra and Biondo Biondi

FIGURES

Figure 1: a arm11phw4 [ER]

Guerra and Biondi  2  Prestack exploding reflector modeling

Figure 5: a arm11wang [ER] SEP–134 Figure 6: a arm11z4ang [ER] SEP–134

Guerra and Biondi  7  Prestack exploding reflector modeling

Figure 7: a armig11T [ER] SEP–134

Guerra and Biondi  13  Prestack exploding reflector modeling

Figure 13: a armig41W [ER] SEP–134 Figure 19: a arr11phw [ER] SEP–134

Guerra and Biondi  24  Prestack exploding reflector modeling

Figure 24: a arrecev [ER] Figure 25: a arrphw [ER] SEP–134 Figure 27: a ars11phw [ER] SEP–134 Figure 28: a ars11phz [ER] SEP–134

Guerra and Biondi  31  Prestack exploding reflector modeling

Figure 32: a arsource [ER] SEP–134 Figure 33: a mig08 [ER] SEP–134 Figure 34: a rot06rec [ER] SEP–134 Figure 35: a rot06sou [ER] SEP–134

Guerra and Biondi  45  Prestack exploding reflector modeling

Figure 46: a migar1acig [ER] SEP–134 Figure 47: a migar1scig [ER] SEP–134 Figure 48: a migsp1acig [ER] SEP–134 Figure 49: a migsp1scig [ER] SEP–134 Figure 50: a migar1rec [ER] SEP–134 Figure 51: a migar1sou [ER] SEP–134

Guerra and Biondi  51  Prestack exploding reflector modeling

Figure 52: a dipnorot [ER] SEP–134 Figure 53: a diprot [ER] SEP–134 Figure 54: a dipshtpr [ER] SEP–134 Figure 55: a dipshtprV [ER] SEP–134 Figure 56: a recrot [ER] SEP–134 Figure 57: a sourot [ER] SEP–134 Figure 59: a migar1sciglo [ER] SEP–134 Figure 61: a migsp1aciglo [ER] SEP–134 Figure 62: a migsp1sciglo [ER] SEP–134 Figure 64: a nodipshtpr [ER] SEP–134

Guerra and Biondi  64  Prestack exploding reflector modeling

Figure 65: a 1 [ER] SEP–134 Figure 73: a 1a [ER] REFERENCES Figure 78: a xtalk3 [ER] SEP–134